OpenStudy (pulsified333):

Find the value of x at which the tangent line to the curve y=1/(x+8) passes through the origin.

myininaya (myininaya):

say we find the tangent line at x=a \[y-f(a)=f'(a)(x-a) \\ y=f'(a)(x-a)+f(a)\] can you find the y-intercept of this tangent line?

OpenStudy (pulsified333):

I'm sorry. I am really confused :(

myininaya (myininaya):

do you know where I got y=f'(a)(x-a)+f(a) ?

myininaya (myininaya):

the point slope form of a line is y-f(a)=m(x-a) agree? if we know this line is tangent to y=1/(x+8) at x=a then m=f'(a) y-f(a)=f'(a)(x-a) adding f(a) on both sides gives y=f'(a)(x-a)+f(a)

myininaya (myininaya):

now actually we are given this line goes through point (0,0) so just enter in (0,0) for (x,y)

OpenStudy (pulsified333):

oh

myininaya (myininaya):

and you will get an equation just in terms of a

myininaya (myininaya):

then you solve that equation for a

OpenStudy (pulsified333):

what about f'(a), do i plug in 0?

myininaya (myininaya):

no

myininaya (myininaya):

f'(a) is the slope at x=a

myininaya (myininaya):

you are given f(x)=1/(x+8) you must find f'(x)

OpenStudy (pulsified333):

f'(x)=-1/(x+8)^2

myininaya (myininaya):

right and that evaluated at a is -1/(a+8)^2

OpenStudy (pulsified333):

right

myininaya (myininaya):

\[y=\frac{-1}{(a+8)^2}(x-a)+f(a)\] you still know this line goes through (0,0) so you still need to enter in (0,0) for (x,y) and I bet you know that f(a)=1/(a+8)

myininaya (myininaya):

\[0=\frac{-1}{(a+8)^2}(0-a)+\frac{1}{a+8}\] solve for a

OpenStudy (pulsified333):

a=-4?

myininaya (myininaya):

looks good

OpenStudy (pulsified333):

that makes so much more sense now :D, thank you!

myininaya (myininaya):

you can check your work by pluggin back -4 for a into y=f'(a)(x-a)+f(a) \[y=\frac{-1}{(-4+8)^2}(x-(-4))+\frac{1}{-4+8} \\ y=\frac{-1}{4^2}(x+4)+\frac{1}{4} \\ y=\frac{-1}{4^2}x-\frac{1}{4}+\frac{1}{4} \\ y=\frac{-1}{4^2}x\] yep this line definitely passes through (0,0)

myininaya (myininaya):

y=mx+b if b is 0 then (0,0) is on the line y=mx+b